OpenStudy (anonymous):

Solve the equation by completing the square: x2 + 2x = 3

OpenStudy (anonymous):

x^2 + 2x + 1/4 = 3 + 1/4

OpenStudy (anonymous):

X²=√X²,X=1 i got this as an answer

OpenStudy (anonymous):

i think im completing it right hang on

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

wait plus 1

OpenStudy (anonymous):

x^2 + 2x (+1) = 3 (+1)

OpenStudy (anonymous):

then you'll get (x+1)^2 = 4 and that's your answer

OpenStudy (anonymous):

i need to find what the two x"s are

OpenStudy (anonymous):

its 1 or -3

OpenStudy (anonymous):

since (1+1)^2 = 4 and (-3+1)^2 both equal 4

OpenStudy (anonymous):

okay i understand now thank you(:

OpenStudy (anonymous):

yep np

OpenStudy (anonymous):

to find it usually you'd sqrt both sides just remember the answer can be pos or neg and solve

OpenStudy (anonymous):

WAITTTTTT

OpenStudy (anonymous):

?

OpenStudy (anonymous):

12 Answer: Solve the equation by completing the square: x2 + 8x = 14

OpenStudy (anonymous):

x^2 + (8x) + ? = 14 + ?

OpenStudy (anonymous):

you are very smart @funkeemonk3y

OpenStudy (anonymous):

then x^2 + 8x + 16 = 14 + 16

OpenStudy (anonymous):

then you'd have (x+4)^2 = 30

OpenStudy (anonymous):

were do you get 16 from?

OpenStudy (anonymous):

note for completing the square its generally a^2x + bx + c to be a complete square c = (b/2)^2

OpenStudy (anonymous):

in order to compete the square thats the main concept you must make it into the form (x + a)^2 = #

OpenStudy (anonymous):

note that (a+b)^2 = a^2 + 2ab + b^2

OpenStudy (anonymous):

gotta go hope this helps

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

x^2+2x+1/4=3+1/4